OpenStudy (anonymous):

A piling for a high-rise building is pushed by two bulldozers at exactly the same time. One bulldozer exerts a force of 1250 pounds in a westerly direction. The other bulldozer pushes the piling with a force of 2650 pounds in a northerly direction. What is the direction of the resulting force upon the piling, to the nearest degree?

OpenStudy (anonymous):

Draw or imagine a right triangle with the vertical side (North) a = 2650 lb Make the bottom side go to the right (East) b = 1250 lb. The hypotenuse will be c^2 = a^2 + b^2, and the total force will be the square root of this. The direction will be such that tan@ = 2650/1250 so you can find @. I should go to bed!